# A New Method to Simulate Dark Matter–Baryon Interactions and Application to an Isolated Disk Galaxy

### 25 Mar 2026 v1

@@ Abstract
We report on a new method for incorporating interactions between dark matter (DM) and baryons in cosmological simulations, capable of handling the challenging regime in which the dark matter particle mass is comparable to or lighter than the baryon mass. The method hybridizes two distinct approaches, as gas particles receive momentum and energy transfer according to a mean-field calculation while DM particles undergo Monte Carlo scatterings, which are derived from the Boltzmann equation and proved to be statistically equivalent. We present an open-source implementation of this method in the simulation code GIZMO. As a first application, we investigate the effects of DM–baryon interactions on an isolated Milky Way-like disk galaxy for dark matter having twice the proton mass, which roughly maximizes the average energy transfer per collision. For cross sections of order 1 barn ($10^{-24}\,\mathrm{cm}^{2}$), these interactions cause strong changes to the mass distribution in the center of the galaxy in less than 1 Gyr, even when bar formation is suppressed by hand. For cross sections typical of hadronic interactions, $\lesssim 30\,{\rm mb}$, high-fidelity galaxy formation simulations will be needed to assess the effects on observable features of galaxies.
@@ I Introduction
The most commonly known particle physics model featuring non-gravitational interactions between DM and the Standard Model is the Weakly Interacting Massive Particle (WIMP). WIMP dark matter is a natural prediction of supersymmetry, but the parameter space for WIMP dark matter is now severely constrained by accelerator limits and direct detection experiments such as Aprile et al. (2023); Aalbers et al. (2025). Moreover, due to its heavy mass and small cross section with baryons, WIMP dark matter is effectively non-interacting except gravitationally.
Here, we are interested in dark matter which has observationally impactful interactions with baryons. Direct detection of DM which is light and has moderate or strong interactions with baryons is challenging. Unless it is weakly interacting, a DM particle loses most of its energy interacting in the material above the detector, leaving too little energy to be detected. Figure 8 of (Albakry and others, 2023) summarizes exclusion regions for cross sections below 1 mb ($10^{-27}\,{\rm cm}^{2}$). Xu and Farrar (2023) reviews the cross section constraints on DM in the GeV mass regime, finding that the direct detection constraints are non-existent or tentative for dark matter masses $\lesssim 3$ GeV and dark matter–baryon interaction cross sections in the $10^{-28}\text{--}10^{-25}\,{\rm cm}^{2}$ range. A proposed limit in this range, based on detection in a deep underground experiment of light DM up-scattered by cosmic rays, will potentially become constraining after non-resonant energy losses in propagation through the Earth are accounted for (Alvey et al., 2023). Limits are also obtained from astrophysical systems such as gas-rich dwarf galaxies, cf., Wadekar and Farrar (2021), clusters, and many others.
A complementary approach to constraining dark matter–baryon (DM–b) interactions is through their astrophysical impacts on the structure of galaxies (e.g., Dvorkin et al., 2014; Buen-Abad et al., 2022). The effects on the structure of galaxies come about in two primary ways: by introducing viscosity before recombination, suppressing small scale power, and by changing the formation and evolution of galaxies. The pre-recombination imprints have been used to place limits on interaction parameters by comparing against measurements of the CMB (Gluscevic and Boddy, 2018; Boddy and Gluscevic, 2018; Boddy et al., 2018; Ali-Haïmoud et al., 2024), the Lyman-$\alpha$ forest (Dvorkin et al., 2014; Xu et al., 2018), large-scale structure (He et al., 2023, 2025), and the abundance of Milky Way satellites (Nadler et al., 2021; Maamari et al., 2021; Buen-Abad et al., 2022).
@@
The purpose of our work is to develop tools to enable an accurate treatment of the effects of dark matter–baryon interactions during galaxy formation and evolution. We have developed a new method for incorporating DM–b interactions into hydrodynamic cosmological simulations, which we have implemented as a new module in the code GIZMO (Hopkins, 2015). The present implementation treats elastic scatterings between non-relativistic dark matter and baryon particles with a phenomenological, isotropic power-law cross section $\sigma(v)=\sigma_{n}\,v^{n}$, with $v\ll c$ the relative velocity, however the method applies for general cross section dependencies on $v$ and scattering angle. The method hybridizes two statistically equivalent but qualitatively different approaches: the dark matter is treated in a Monte Carlo fashion, where simulation particles probabilistically undergo scatterings, and the baryons are treated with a mean field approach, where the statistical expectations for the transfer rates of bulk velocity and thermal energy are applied. These are both derived directly from the Boltzmann equation, ensuring that energy and momentum transfer between the sectors are equal and opposite on average.
During the preparation of this article, we learned of a parallel effort to implement dark matter–baryon interactions in $N$-body simulation from Fischer et al. (2025). There, the simulation dark matter particles scatter off of “virtual” baryon particles, with the resulting energy and momentum change of the virtual particle distributed back onto the actual simulation baryon macroparticles. While we defer a more detailed comparison with their method to the main text, we note that there are two important differences between our methods. The Fischer et al. (2025) approach conserves energy and momentum exactly per scattering, while we give this up and conserve energy and momentum only on average. However, their virtual particle scheme requires that the masses of the simulation macroparticles are in a similar ratio as the masses of the elementary dark matter to protons. In high-quality simulations of galaxies, much less resolution is needed for the DM particles than for baryons without DM–b interactions, and DM macroparticles masses can be orders of magnitude heavier than those for baryons. Thus to apply the F25 method with comparable mass DM and baryon macroparticles may be computationally infeasible for $m_{\chi}$ comparable to or less than the mass of the proton. By contrast, our approach—by treating the baryons separately from the probabilistic DM–b scattering—sidesteps this issue entirely, enabling us to press into light dark matter masses without huge computational overhead.
@@
Our method thus provides for the first time a practical tool designed for investigating dark matter–baryon interactions with dark matter mass close to that of hydrogen and helium, where energy and momentum transfer in individual collisions is maximally efficient. A concrete example of such a model is Sexaquark Dark Matter, in which the dark matter particle is a six-quark state (uuddssuuddss) whose binding is sufficiently deep that it is effectively stable; see Farrar (2022) and references therein for a review of the constraints from particle physics, astrophysics, and cosmology.
@@
As an initial application of our code, we simulate the isolated Milky Way-like disk galaxy of the AGORA project (Kim et al., 2016) with a dark matter particle having twice the proton mass ($2\,m_{p}$) and velocity-independent cross section. We find that sufficiently strong dark matter–baryon interactions induce significant changes to the morphology of the center of the galaxy. In particular, such interactions drive gas and dark matter inflows toward the center. For potentially realistic DM–baryon cross sections, below tens of millibarns ($10\ {\rm mb}=10^{-26}\ {\rm cm}^{2}$), assessing the impact of dark matter–baryon interactions will require more detailed treatment of star formation and feedback, and of the hierarchical formation of galaxies, than we undertake in the present work.
@@ II.1 Boltzmann equation
where the dependence on $\bm{x}$ is implicit. The right-hand-side of this equation is known as the collision integral.
@@
The first approximation made by all $N$-body simulations is that the dark matter is a “softened $N$-body component”; namely, that a number of Monte Carlo samples are drawn from some true initial df $f_{\chi}$ and spatially convolved with a smoothing kernel W(r)W(r), [^2] We assume throughout that any valid smoothing kernel $W(r,h)$ integrates to 1 and is zero (or else negligible) for $r>h$. producing the following approximate df
$\hat{f}_{\chi}(\bm{x},\bm{v})=\sum_{i}\frac{M_{i}}{m_{\chi}}\,W(|\bm{x}-\bm{x}_{i}|,h_{i})\,\delta(\bm{v}-\bm{v}_{i}).$ (4)
@@
The second approximation is that the gas is represented by a thermalized softened $N$-body component. By this, we mean that the gas is decomposed into macroparticles like dark matter, but each macroparticle has a bulk velocity and internal energy describing the local velocity distribution, which is taken to be Maxwell-Boltzmann at all times. Therefore, the gas is represented by the approximate df
$\hat{f}_{B}(\bm{x},\bm{v})=\sum_{j}\frac{M_{j}}{\mu_{j}}\,W(|\bm{x}-\bm{x}_{j}|,h_{j})\,\mathcal{G}(\bm{v}-\bm{V}_{j},T_{j}/\mu_{j}),$ (5)
where $\mathcal{G}$ is the probability density function of a three-dimensional isotropic Gaussian—i.e., the Maxwell-Boltzmann distribution. In simulation, then, the gas is represented by $N_{B}$ macroparticles, each having a mass $M_{j}$, position $\bm{x}_{j}$, smoothing length $h_{j}$, bulk velocity $\bm{V}_{j}$, temperature $T_{j}$, and mean molecular weight $\mu_{j}$.
We note that the form of the baryon df given in Equation 5 is valid regardless of the choice of hydrodynamic solver. By computing the moments of the Boltzmann equation for exactly this df, one can derive the fluid equations for standard smoothed particle hydrodynamics (e.g., Monaghan, 1992). More modern solvers, such as the moving Voronoi mesh of AREPO (Springel, 2010) or the meshless methods of GIZMO (Hopkins, 2015), generally assume sophisticated geometries for discretizing the gas. However, in these simulations the resulting particle distribution is then taken to have the form of Equation 5 for other applications, such as solving gravity, and so we take this form for solving interactions with dark matter.
@@
The third assumption underpinning these simulations is that the time-evolution of the true dfs $f_{\chi},f_{B}$ is well-approximated even when enforcing that the form of the macroparticle dfs $\hat{f}_{\chi},\hat{f}_{B}$ is preserved. In other words, after a timestep, the forms of Equations 4 and 5 can be imposed on the dark matter and gas, albeit with updated values of the parameters. We make this point explicit because the full evolution of the df according to the Boltzmann equation would distort the functional form of the discretization—for example, the softened macroparticles would experience tidal forces and become stretched and sheared—but this is regularly ignored. For our purposes, this assumption means we can impose that evolution according to the collision integral does not change the form of the discretization: after an update, the dark matter will have the form Equation 4 and the baryons will have the form Equation 5. Because the collision integral is local, it will also not change the masses or positions of any particles. Thus, we need only consider the update to the individual particle velocity distributions, and so the Boltzmann equation is
@@
and similarly for the baryons $\hat{f}_{B}$.
The form of this equation suggests a matching procedure: term $i$ on the left-hand-side can be set equal to term $i$ on the right-hand-side. By doing this matching, we are making an additional assumption: that the phase space densities of macroparticles of the same type do not mix, so each macroparticle evolves as if it is the only particle of its type at its location. This assumption is reasonable as long as no particles have identical positions $\bm{x}$, which is already required for the validity of the $N$-body approach. Hence, individual macroparticles evolve according to independent collision integrals:
@@ II.2 Dark matter evolution
$f_{\chi}^{i}\bigg|_{t+\Delta t}=f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t$ (9)
@@
The derivation proceeds from here in detail in Appendix A. Briefly, the steps we take are as follows. We substitute into Equation 9 the baryon macroparticle df $\hat{f}_{B}$. This gives a sum over gas particles $j$, each with a different integral over ${\rm d}^{3}\bm{v}_{B}\,{\rm d}\Omega$. We simulate this integral by drawing a single Monte Carlo sample independently for every $j$ (denoting the sampled baryon velocity $\bm{v}^{B}_{ij}$ and scattering angles $\Omega_{ij}$). Then, in order to isolate the velocity distribution on both sides, we integrate both sides over all space. The result is
@@
and $v_{\rm rel}\equiv|\bm{v}_{i}-\bm{v}_{ij}^{B}|$. Note the appearance of the total cross section $\sigma(v)$. We have introduced two intermediate variables to clarify the meaning of this result. One is $g_{ij}$, introduced by Rocha et al. (2013) for self-interacting dark matter simulations and dubbed the “geometric factor” or the “overlap integral”. It has dimensions of inverse volume and gives the effective number density of particle $j$ acting on particle $i$. The other is $P_{ij}$ which is dimensionless and, as long as $P_{ij}\ll 1$, can be understood as a probability. Indeed, $(M_{j}/\mu_{j})\,g_{ij}$ is the effective number density of microscopic baryons active in this interaction, so $P_{ij}$ has the form $n\,\sigma\,v\,\Delta t$, exactly as one would expect.
The new velocity distribution $\nu_{i}^{\prime}$ is a discrete probability distribution: it gives the post-scattering velocity $\bm{S}_{\chi}(\bm{v}_{i},\bm{v}^{B}_{ij},\Omega_{ij})$ with probability $P_{ij}$, or the original velocity $\bm{v}_{i}$ with probability $1-\sum_{j}P_{ij}$. To pick a new velocity for macroparticle $i$, then, we can draw a sample from this discrete distribution. Note that the time-evolution of the dark matter macroparticles is given by probabilistically scattering them off of samples from the baryon velocity distribution in a manner which agrees exactly with the microscopic theory.

### 17 Apr 2026 v2

@@ Abstract
We report on a new method for incorporating interactions between dark matter (DM) and baryons in cosmological simulations, capable of handling the challenging regime in which the dark matter particle mass is comparable to or lighter than the baryon mass. The method hybridizes two distinct approaches: gas particles receive momentum and energy transfer according to a mean-field calculation while DM particles undergo Monte Carlo scatterings. These approaches are derived from the Boltzmann equation and shown to be statistically equivalent. We present an open-source implementation of this method in the simulation code GIZMO. As a first application, we investigate the effects of DM–baryon interactions on an isolated Milky Way-like disk galaxy for dark matter having twice the proton mass, which roughly maximizes the average energy transfer per collision. For cross sections of order 1 barn ($10^{-24}\ {\rm cm}^{2}$), these interactions cause strong changes to the mass distribution in the center of the galaxy in less than 1 Gyr, even when bar formation is suppressed by hand. For cross sections typical of hadronic interactions ($\lesssim 30\ {\rm mb}$), high-fidelity galaxy formation simulations will be needed to assess the effects on observable features of galaxies.
@@ I Introduction
Dark matter–baryon (DM–b) interactions are a feature of many particle models of dark matter. The most common such model is the Weakly Interacting Massive Particle (WIMP), although it is not very interesting observationally due to its heavy mass, small cross section with baryons, and the severe constraints imposed by accelerator limits and direct detection experiments (e.g., Aprile et al., 2023; Aalbers et al., 2025). We focus on dark matter with more observationally impactful interactions with baryons, such as the regime of light mass DM with moderate-to-strong interactions with baryons that is challenging to directly detect. In the GeV mass range, deep underground direct detection experiments like SuperCDMS are sensitive only to weak interactions (cross sections of $10^{-35}$ to $10^{-29}\ {\rm cm}^{2}$) (Albakry and others, 2023), leaving the space $\gtrsim 10^{-28}\ {\rm cm}^{2}$ nearly unconstrained (Xu and Farrar, 2023). [^1] It has been proposed that the upscattering of DM by cosmic rays can enhance the sensitivity of direct detection experiments (Alvey et al., 2023), which could significantly constrain the parameter space in this range after non-resonant energy losses in propagation through the Earth are accounted for.
A complementary approach to constraining DM–b interactions is through their astrophysical impacts. For example, Wadekar and Farrar (2021) placed limits by requiring that the heating or cooling due to DM–b interactions not exceed the radiative cooling rate of gas in gas-rich dwarf galaxies. DM–b interactions can also affect the structure of galaxies (e.g., Dvorkin et al., 2014; Buen-Abad et al., 2022), primarily in two ways: by introducing viscosity before recombination and thereby suppressing the power spectrum on small scales, and by changing the formation and evolution of galaxies. The pre-recombination imprints have been used to place limits on interaction parameters by comparing against measurements of the CMB (Gluscevic and Boddy, 2018; Boddy and Gluscevic, 2018; Boddy et al., 2018; Ali-Haïmoud et al., 2024), the Lyman-$\alpha$ forest (Dvorkin et al., 2014; Xu et al., 2018), large-scale structure (He et al., 2023, 2025), and the abundance of Milky Way satellites (Nadler et al., 2021; Maamari et al., 2021; Buen-Abad et al., 2022).
@@
The purpose of our work is to develop tools to enable an accurate treatment of the effects of dark matter–baryon interactions during galaxy formation and evolution. We have developed a new method for incorporating DM–b interactions into hydrodynamic cosmological simulations, which we have implemented as a module in the code GIZMO (Hopkins, 2015). We focus on elastic scatterings between non-relativistic dark matter and baryon particles with a phenomenological, isotropic power-law cross section $\sigma(v)=\sigma_{0}\,(v/c)^{n}$, where $v\ll c$ is the relative velocity, though we note that our methods can be applied to general cross section dependencies on $v$ and scattering angle. The method hybridizes two statistically equivalent but qualitatively different approaches: the dark matter is treated in a Monte Carlo fashion, where simulation particles probabilistically undergo scatterings, and the baryons are treated with a mean-field approach, where the statistical expectations for the transfer rates of bulk velocity and thermal energy are applied. These are both derived directly from the Boltzmann equation and proven to conserve energy and momentum in expectation.
During the preparation of this article, we learned of a parallel effort to implement DM–b interactions in $N$-body simulations (Fischer et al., 2025) (F25 below). The advantages and disadvantages of that approach relative to ours are discussed in Section II.5. The most significant difference is that the F25 approach requires the dark matter and baryon macroparticle masses to satisfy $M_{\chi}/M_{b}<3/28\,(m_{\chi}/m_{b})\approx 0.1\,m_{\chi}/{\rm GeV}$, so employing the F25 method in the $m_{\chi}\lesssim$ few GeV regime requires a much larger number of dark matter macroparticles than for non-interacting DM. [^2] For reference, state-of-the-art cosmological simulations adopt $M_{\chi}/M_{b}\approx 5$ (e.g., Hopkins et al., 2018; Nelson et al., 2019, 2024), and some isolated galaxy sims choose ratios as high as $M_{\chi}/M_{b}\approx 100$ (Kim et al., 2016).
@@
Our method thus provides for the first time a practical tool for investigating dark matter–baryon interactions for a dark matter mass close to that of hydrogen and helium, where energy and momentum transfer in individual collisions is maximally efficient. A concrete example of such a model is Sexaquark Dark Matter, in which the dark matter particle is a six-quark state (uuddssuuddss) whose binding is sufficiently deep that it is effectively stable; see Farrar (2022) and references therein for a review of the constraints from particle physics, astrophysics, and cosmology, and Moore and Profumo (2025) for additional discussion.
@@
As an initial application of our code, we simulate the isolated Milky Way-like disk galaxy of the AGORA project (Kim et al., 2016) with a dark matter particle having twice the proton mass ($2\,m_{p}$) and velocity-independent cross section. We find that sufficiently strong dark matter–baryon interactions induce significant changes to the morphology of the center of the galaxy, driving gas and dark matter inflows toward the center. For potentially realistic DM–b cross sections below tens of millibarns, assessing the impact of dark matter–baryon interactions will require more detailed treatment of star formation, feedback, and the hierarchical formation of galaxies than we undertake in the present work.
@@ II Method
In this section, we derive our method and discuss our implementation. We begin by establishing the Boltzmann equation of our simulations using three core assumptions that underpin $N$-body simulations. We use this Boltzmann equation to derive both a probabilistic scattering treatment for dark matter, and the mean acceleration and heating rates for baryons. Then, we describe our implementation of this method in the simulation code GIZMO, and finally we provide a detailed comparison of our method with that of Fischer et al. (2025).
@@ II.1 Boltzmann equation
where the dependence on $\bm{x}$ is implicit. The right-hand side of this equation is known as the collision integral.
@@
The first approximation made by all $N$-body simulations is that the dark matter is a “softened $N$-body component”; namely, a number of Monte Carlo samples are drawn from some true initial df $f_{\chi}$ and spatially convolved with a smoothing kernel W(r)W(r), [^4] We assume throughout that any valid smoothing kernel $W(r,h)$ integrates to 1 and is zero (or else negligible) for $r>h$. producing the following approximate df
$\hat{f}_{\chi}(\bm{x},\bm{v})=\sum_{i=1}^{N_{\chi}}\frac{M_{i}}{m_{\chi}}\,W(|\bm{x}-\bm{x}_{i}|,h_{i})\,\delta(\bm{v}-\bm{v}_{i}).$ (4)
@@
The second approximation is that the gas is represented by a thermalized softened $N$-body component. By this, we mean that the gas is decomposed into macroparticles like dark matter, but each macroparticle has a bulk velocity and internal energy describing the velocity distribution, which is locally Maxwell-Boltzmann at all times. Therefore, the gas is represented by the approximate df [^5] We note that this form of the baryon df is valid for any hydrodynamic solver. By computing the moments of the baryon–baryon Boltzmann equation for this df, one can derive the fluid equations for standard smoothed particle hydrodynamics (e.g., Monaghan, 1992). More modern solvers, such as the moving Voronoi mesh of AREPO (Springel, 2010) or the meshless methods of GIZMO (Hopkins, 2015), generally assume more sophisticated geometries for the discretization of the gas. However, in these simulations the resulting particle distribution is then taken to have the form of Equation 5 for other applications, such as solving gravity, and so we take this form for solving interactions with dark matter.
$\hat{f}_{B}(\bm{x},\bm{v})=\sum_{j=1}^{N_{B}}\frac{M_{j}}{\mu_{j}}\,W(|\bm{x}-\bm{x}_{j}|,h_{j})\,\mathcal{G}(\bm{v}-\bm{V}_{j},T_{j}/\mu_{j}),$ (5)
where $\mathcal{G}$ is the probability density function of a three-dimensional isotropic Gaussian—i.e., the Maxwell-Boltzmann distribution. In simulation, the gas is represented by $N_{B}$ macroparticles, each having a mass $M_{j}$, position $\bm{x}_{j}$, smoothing length $h_{j}$, bulk velocity $\bm{V}_{j}$, and temperature $T_{j}$. Gas macroparticles can also consist of different species of particles; we approximate the dark matter as having equal cross section with all species, so all that matters is the mean molecular weight, $\mu_{j}$.
@@
The third approximation underpinning these simulations is made by assuming that the time-evolution of the true dfs $f_{\chi},f_{B}$ is well-approximated even when enforcing that the form of the macroparticle dfs $\hat{f}_{\chi},\hat{f}_{B}$ is preserved. In other words, after a timestep, the forms of Equations 4 and 5 can be imposed on the dark matter and gas, albeit with updated values of the parameters. We make this point explicit because the full evolution of the df according to the Boltzmann equation would distort the functional form of the discretization—for example, the softened macroparticles would experience tidal forces and become stretched and sheared—but this is regularly ignored. For our purposes, this assumption means that evolution according to the collision integral only updates the local velocity distributions of the individual particles. Thus, the Boltzmann equation is
@@
and similarly for the baryons.
The form of this equation suggests a matching procedure: term $i$ on the left-hand side can be set equal to term $i$ on the right-hand side. By doing this matching, we are making an additional assumption: that the phase space densities of macroparticles of the same type do not mix, so each macroparticle evolves as if it is the only particle of its type at its location. This assumption is reasonable as long as no particles have identical positions $\bm{x}$, which is already required for the validity of the $N$-body approach. Hence, individual macroparticles evolve according to independent collision integrals:
@@ II.2 Dark matter evolution
$f_{\chi}^{i}\bigg|_{t+\Delta t}=f_{\chi}^{i}\bigg|_{t}+\frac{{\rm d}f_{\chi}^{i}}{{\rm d}t}\bigg|_{t}\,\Delta t,\vskip 5.0pt$ (9)
@@
The derivation proceeds from here in detail in Appendix A. Briefly, the steps we take are as follows. We substitute into Equation 9 the baryon macroparticle df $\hat{f}_{B}$. This gives a sum over gas particles $j$, each with a different integral over ${\rm d}^{3}v_{B}\,{\rm d}\Omega$. We simulate this integral by drawing a single Monte Carlo sample independently for every $j$ (denoting the sampled baryon velocity $\bm{v}^{B}_{ij}$ and scattering angles $\Omega_{ij}$). Then, in order to isolate the velocity distribution on both sides, we integrate both sides over all space. The result is
@@
and $v_{\rm rel}\equiv|\bm{v}_{i}-\bm{v}_{ij}^{B}|$. Note the appearance of the total cross section $\sigma(v)$. We have introduced two intermediate variables to clarify the meaning of this result. One is $g_{ij}$, introduced by Rocha et al. (2013) for self-interacting dark matter simulations and dubbed the “geometric factor” or the “overlap integral”. It has dimensions of inverse volume and gives the effective number density of particle $j$ acting on particle $i$. The other is $P_{ij}$, which is dimensionless and, as long as $P_{ij}\ll 1$, can be understood as a probability. Indeed, $(M_{j}/\mu_{j})\,g_{ij}$ is the effective number density of microscopic baryons active in this interaction, so $P_{ij}$ has the form $n\,\sigma\,v\,\Delta t$, exactly as one would expect.
The new velocity distribution $\nu_{i}^{\prime}$ is a discrete probability distribution: it gives the post-scattering velocity $\bm{S}_{\chi}(\bm{v}_{i},\bm{v}^{B}_{ij},\Omega_{ij})$ with probability $P_{ij}$, or the original velocity $\bm{v}_{i}$ with probability $1-\sum_{j}P_{ij}$. To pick a new velocity for macroparticle $i$, we can draw a sample from this discrete distribution. Notice that the time-evolution of the dark matter macroparticles is given by probabilistically scattering them off samples from the baryon velocity distribution in a manner that agrees exactly with the microscopic theory.